• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name E. Schulte

    Log entry time 15:58:27 on November 05, 2008

    Entry number 246925

    keyword=Day Shift Summary, November 5, 2008

    08:00 Shift starts taking data on the reference cell with hydrogen
    pressurized to 110 psi. There exists some confusion about the plan for
    the next hour or so. But, we're trying to sort it out.

    08:30 We got it all straightened out. We're going to run in the current
    configuration (reference cell with hydrogen, BigBite Sieve in, target
    collimator out) until further notice. At the time of further notice
    (which will conveniently coincide with a beam down period) we will make
    the access to remove the BigBite sieve, and put in the target collimator.

    09:00 Doug calls for Brad, something about the beam...Need to make sure
    Brad calls him back.

    09:06 Vince & Kalyan need to do some runs to locate the coincidence time.

    09:38 MCC called to find out if we had the dummy target in. I told them
    it's the reference cell.

    10:00 MCC called to let us know that they have an IOC issue and will be
    down for an undetermined amount of time. Called for a Controlled Access.

    10:05 MCC is also going to do a Wien angle change. And the target
    polarization is 40%!!!

    10:46 Put Reference Cell Control Panel in Local Mode.

    11:22 Collimators are an and Yi is going to unlock the lasers.

    11:46 Reference Cell Control Panel now back in Remote Mode. The
    regulator has been replaced so we can go to 135 psi on the reference cell
    with hydrogen. The current pressure is 100 psi.

    11:49 RC lets us know that MCC is still having issues with an IOC, so
    we're still on "their time"!

    12:25 Called MCC to see what's up. They're still working on their IOC
    and will call us when they're done.

    12:35 MCC ready to send RF to hall. Request Beam Permit.

    12:52 Request tune beam (2uA) from MCC to check the positions.

    13:00 Requested change of target to Optics (BeO).

    13:03 Requested 2uA of tune back to check spot on BeO target.

    13:23 Request a harp scan from MCC. Changed target to "empty" position.

    13:33 Harp scan complete. Harp B scan looks pretty crummy. D. Gaskell
    kindly logged them.

    13:39 Requested 5uA with MCC 4x4 raster. Check with spot++. Then the
    beam got really unstable.

    13:54 Change to optics target.

    13:55 Optics target run with raster of, 5uA. Target collimator in, Sieve
    in. The beam has been quite unstable.

    14:10 Request 4x4 raster from MCC.

    14:12 Begin filling reference cell with hydrogen to 135 psi. Started at
    100 psi. The fill was successful.

    14:26 Start raster scan. Reference cell, hydrogen 135 psi

    Hall A raster values & T1 Rates
    2x2 2 kHz
    3x3 2 kHz
    4x4 2 kHz
    5x5 2 kHz

    14:33 Request 4x4 raster, 5uA on reference cell with 135 psi hydrogen.
    We will run for 2 hours beam-on-target.

    14:38 CODA seems to have died. We'll try rebooting it. ROC's 8 & 9,
    which died.

    14:47 ROC's recovered, restarted taking data.

    14:53 Gas Shed Airflow Alarm (cleared itself)

    15:15 Merrily taking data. Beam's been a little better the last half
    hour or so.

    13:57 Shift ends taking data on reference cell, with hydrogen 135 psi.
    We have about an hour or so of beam on target.